OpenStudy (megannicole51):

Find the sum of the series. For what values of the variable does the series converge to this sum? 1+(x/2)+(x^2/4)+(x^3/8)+....

OpenStudy (zarkon):

Looks like \[\sum_{k=0}^{\infty}\left(\frac{x}{2}\right)^k\]

OpenStudy (megannicole51):

okay:)

OpenStudy (zarkon):

This site is driving me crazy with the lag...oh well you need \[\left|\frac{x}{2}\right|<1\] in order to have convergence

OpenStudy (megannicole51):

can i use the infinite or finite equation?

OpenStudy (zarkon):

what do you mean?

OpenStudy (megannicole51):

a/1-r or a(1-r^n/1-r)

OpenStudy (megannicole51):

@agent0smith do u know how to do this?

OpenStudy (megannicole51):

my test is in like 3 hours! its crunch time!!!

OpenStudy (megannicole51):

@Coolsector do u know how to do this?

OpenStudy (anonymous):

so it is a geometric series sum = 1/(1-q)

OpenStudy (anonymous):

where q in our case is x/2

OpenStudy (anonymous):

and it will converge only if |q|<1

OpenStudy (megannicole51):

how do u know that its a geometric series? u just subtract the third term from the second and the second from the first term to see if they are the same....which is also "r" correct?

OpenStudy (megannicole51):

well q in ur case

OpenStudy (agent0smith):

^ divide, remember Megan, subtracting is for an arithmetic series. and if you divide, yes it should be geometric. Geometrics only converge is r < 1.

OpenStudy (megannicole51):

and how do u know to use that formula and not a((1-r^n)/(1-r))?

OpenStudy (anonymous):

yes you see that every term is multiplied by x/2

OpenStudy (anonymous):

a((1-r^n)/(1-r)) - > this is for the finite case

OpenStudy (agent0smith):

^ that formula only works if a geometric series converges.

OpenStudy (agent0smith):

^and that too.

OpenStudy (agent0smith):

btw this question wasn't asking you to find the sum, that's what that formula and the infinite formula are for. It's only asking you to find what values of x is it even possible to get an infinite sum.

OpenStudy (megannicole51):

how can i tell the difference between a finite and infinite series?

OpenStudy (anonymous):

the "..." at the end without specific term means that it is infinite

OpenStudy (dan815):

gotta do some tests

OpenStudy (dan815):

nvm i thought u mean for convergence or diveregence

OpenStudy (anonymous):

if otherwise it was 1+(x/2)+(x^2/4)+(x^3/8)+....+ (x^8/2^8) you would know that is finite

OpenStudy (megannicole51):

beautiful thats exactly what i needed to know...thank you!

OpenStudy (anonymous):

yw

OpenStudy (megannicole51):

so since i am finding the sum of this series and it has ..... at the end its infinite?

OpenStudy (anonymous):

infinite series yes

OpenStudy (megannicole51):

so i use the formula a/1-r?

OpenStudy (anonymous):

yes and the condition for the sum to converge is |r| < 1

OpenStudy (agent0smith):

don't forget you still need to solve this for x (since this is your r) \[\large \left|\frac{x}{2}\right|<1\]since that'll give the values it converges for.

OpenStudy (megannicole51):

beautiful

OpenStudy (megannicole51):

how is x/2 my r? and why do i have to solve for x? @agent0smith

OpenStudy (megannicole51):

wouldnt it be x^n-x/2?

OpenStudy (agent0smith):

(x/2) is your r since if you divide one term by the one before it, or just notice that each term is multiplied by x/2 to get the next one - that is r.

OpenStudy (agent0smith):

This is why you have to solve for x "For what values of the variable does the series converge to this sum?"

OpenStudy (megannicole51):

okay i got it thank you:)))

OpenStudy (agent0smith):

k :)